3- Suppose country’s population grows at the rate of 2% per year and its output grows at the rate of 3% per year.
a. Calculate its rate of growth of per capita output
1
Expert's answer
2020-10-13T07:27:04-0400

Let the original population be "x"

and the original total output be "y"

Then,

the original per capita output will be "\frac{y}{x}"

According to question ,

New population "= x+\frac{2x}{100} = \frac{102x}{100}"

Then,

New Total Output "= y+ \frac{3y}{100} = \frac{103x}{100}"

So,

New per capita output "=\frac{103y}{102x}"

Then,

Change in per capita output "\frac{103y}{102x} - \frac{y}{x} = \frac{103y-102y}{102x} = \frac{y}{102x}"

Percentage change in per capita output "= \frac{\frac{y}{102x}}{\frac{y}{x}} \times 100" "= \frac{100}{102} = 0.98039216\%"